(this question is commodities related)

Assume that you pay a premium of 30 cents a bushel for a Nov Soybean put option with a $12.50 strike price, and the basis is expected to be 25 cents under November futures when you sell your crop in October. What would your selling price be if the Nov Soybean futures price at expiration (i.e., no time value) is the price shown in the left-hand column?

November Futures Net Return

$11.80 $ _______ per bu

$12.60 $ _______ per bu

$14.30 $ _______ per bu
Answer #1

selling price if the Nov Soybean futures price at expiration 11.8=> 12.5-(0.3+0.25) = $11.95
selling price if the Nov Soybean futures price at expiration 12,6=> 12.6-(0.3+0.25) = $12.05
selling price if the Nov Soybean futures price at expiration 14.3=> 14.3-(0.3+0.25) = $13.75